OpenStudy (anonymous):

please help stuck on this Solve for x: |2x + 6| - 4 = 20 Select one: a. x = 9 and x = 11 b. x = -9 and x = 15 c. x = 9 and x = -15 d. No solution

OpenStudy (anonymous):

it would be C

OpenStudy (campbell_st):

add 4 to both sides of the equation and you'll get |2x + 6| = 24 so you are now solving 2 equations 2x + 6 = 24 and 2x + 6 = -24 hope it helps

OpenStudy (campbell_st):

it would be C

OpenStudy (anonymous):

http://www.wolframalpha.com

OpenStudy (anonymous):

thanks

OpenStudy (tkhunny):

Make your life as easy as you can. Solve for x: |2x + 6| - 4 = 20 |2x + 6| = 24 2|x + 3| = 24 |x + 3| = 12 Now, x + 3 = 12 or x + 3 = -12